OpenStudy (anonymous):

which ordered pair is the solution of this equation -2x - 3y=19 a.-5,-4 b. -5,-2 c. -2,-5 d. -3,-5

OpenStudy (anonymous):

plug each 'solution' into the original problem to solve. :: -2(-5) - 3(-4) = 10 + 12 = 12 (does not =) 19 so 'a' is not the solution. -2(-5) - 3(-2) = 10 + 6 = 16 (does not =) 19 so 'b' is not the solution. -2(-2) - 3(-5) = 4 + 15 = 19 so 'c' is the solution. -2(-3) - 3(-5) = 6 + 15 = 21 so 'd' is not the solution. C is your solution.

OpenStudy (anonymous):

thank you